OpenStudy (anonymous):

write 6 1/4% as a decimal?

OpenStudy (anonymous):



OpenStudy (anonymous):

thank you, that helps me understand

OpenStudy (jdoe0001):

\(\bf a\%\ =\cfrac{a}{100}\qquad thus \\ \quad \\ 6\frac{1}{4}\%\implies \cfrac{6\frac{1}{4}}{100}\implies \cfrac{\frac{25}{4}}{100}\implies \cfrac{25}{4}\cdot \cfrac{1}{100}\implies \cfrac{25}{400}\)